OpenStudy (anonymous):

Solve for x: |3x + 2| = 2

jimthompson5910 (jim_thompson5910):

Hint: If |3x + 2| = 2 then 3x + 2 = 2 or 3x + 2 = -2

jimthompson5910 (jim_thompson5910):

So you need to solve both equations to find your answers.

OpenStudy (anonymous):

you've been asking the same type of questions for awhile now, you should just follow the instructions given to you, instead of just trying to get answers..not being mean or rude...just honest